A 48.3-mL sample of gas in a cylinder is warmed from 24 ∘C to 89 ∘C.

What is its volume at the final temperature? (Assume constant pressure.)
Answer #1

V1 = 48.3 mL

T1 = 24 oC = 297.15 K

T2 = 89 oC = 362.15 K

V2 = final volume

V1 / T1 = V2 / T2

Final volume (V2) = (V1T2) / T1

V2 = (48.3 mL x 362.15 K) / (297.15 K) = 58.86 mL